OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? Answer 7 adults and 4 kids 8 adults and 2 kids 3 adults and 12 kids 5 adults and 6 kids

OpenStudy (anonymous):

plug in each to see which DOES not equal 72

OpenStudy (mertsj):

8(5)+4(6)=40+24=64 Doesn't seem like that one works since it does not equal 72

OpenStudy (anonymous):

8(7) +4(4) = 56+16 =72 ok 64 + 8 = 72 ok 24 + 48 = 72 ok 40 + 24 = 64 ****** your answer

OpenStudy (anonymous):

the answer is 7 adults 4 kids 7x8= 56 4x4=16 56+16=72